What would be the word expression for c= 15,20,25
Answer by jim_thompson5910(35256)

Those look like the third, fourth, and fifth positive multiples of 5.

-------------------------------------------------------

Side Note:

First positive multiple of 5: 1*5 = 5

Second positive multiple of 5: 2*5 = 10

Third positive multiple of 5: 3*5 = 15

Fourth positive multiple of 5: 4*5 = 20

Fifth positive multiple of 5: 5*5 = 25

And so on...